You want to make 500 mL of 0.400 M sodium acetate buffer pH 4.9. You have 17.5 M acetic acid and sodium acetate. Add 5.55 g of sodium acetate. How much acetic acid to add?
Answer #1

Sol.

Let moles of acetic acid be x and moles of sodium acetate be y  

pKa of acetic acid = 4.74

pH of buffer = 4.9

Using Henderson - Hasselbalch equation ,

pH = pKa + log ( moles of sodium acetate / moles of acetic acid )

4.9 = 4.74 + log ( y / x )

log ( y / x ) = 0.16

y / x = 100.16 = 1.4454

Now , mass of sodium acetate = 5.55 g

molar mass of sodium acetate = 82.0343 g / mol

So , moles of sodium acetate = y = 5.55 / 82.0343 = 0.0676 mol

Therefore , moles of acetic acid = x =

y / 1.4454 = 0.0676 / 1.4454 = 0.0468 mol

Conc. of acetic acid = 17.5 M

So , Volume of acetic acid used

= moles of acetic acid / Conc. of acetic acid

= 0.0468 / 17.5

= 0.00267 L

=   2.67 mL